# PACIFIC: a framework for generating benchmarks to check Precise Automatically Checked Instruction Following In Code

### v1

@@ 3. PACIFIC’s Design
Figure 1. PACIFIC: High-level Architecture. Instruction Pool and Benchmark generation parameters feed the Benchmark Generator, producing a Benchmark. Inference runs on the benchmark, and the Evaluation Engine outputs metrics in the form of a report file.:
@@ 3.2. Benchmark Generation Process
Figure 2. Benchmark Sample Structure. A sample consists of an initial input followed by a sequence of instructions forming the prompt, which is processed by the model to generate the output.:
@@
3.3. How PACIFIC Meets the Criteria
@@ 5.1.1. Performance Across Difficulty Levels
Figure 3. Prompt Level Accuracy across varying complexity levels. Higher instruction counts and output lengths correspond to increased task difficulty. Average across all programming languages.:
@@ 5.1.3. Model Comparison
Figure 4. Aggregate performance of evaluated LLMs, measured as the mean Prompt Level Accuracy across all benchmarks.:
@@
Figure 5. Aggregate missing answers rates across models and instruction counts. The higher the instruction count, the higher the error rate.:
@@ 5.1.4. Instruction Samples
Figure 6. Accuracy across models for five representative instructions, grouped by complexity. The to_roman and next_perfect_squared are considered easy, the other ones are the hardest in our current instruction set. Scores are aggregated by instruction count and targeted output length.:
@@ A.1. Current Set of Natural Language Instructions
    Take the previous string - let’s call it s, and output the concatenation of s with itself two times s+s+s.
@@ A.3. Per-Model Performance Details
Figure 7. Claude-3.5-Haiku Performance. Results across all benchmarks for Metric Prompt Level Accuracy.:
@@ A.3.2. Claude-4-Sonnet
Figure 8. Claude-4-Sonnet Performance. Results across all benchmarks for Metric Prompt Level Accuracy.:
@@ A.3.3. GPT-OSS-120B
Figure 9. GPT-OSS-120B Performance. Results across all benchmarks for Metric Prompt Level Accuracy.:
@@ A.3.4. Llama-4-Maverick-17B-128E-Instruct
Figure 10. Llama-4-Maverick-17B-128E-Instruct Performance. Results across all benchmarks for Metric Prompt Level Accuracy.:
@@ A.3.5. Qwen3-235B-A22B-Instruct
Figure 11. Qwen3-235B-A22B-Instruct Performance. Results across all benchmarks for Metric Prompt Level Accuracy.:
@@ A.3.6. Qwen3-Coder-480B-A35B-Instruct
Figure 12. Qwen3-Coder-480B-A35B-Instruct Performance. Results across all benchmarks for Metric Prompt Level Accuracy.:

### v2

@@ 3. PACIFIC’s Design
Figure 1. PACIFIC: High-level Architecture. Instruction Pool and Benchmark generation parameters feed the Benchmark Generator, producing a Benchmark. Inference runs on the benchmark, and the Evaluation Engine outputs metrics in the form of a report file.
PACIFIC high-level architecture diagram showing the main system components.
@@ 3.2. Benchmark Generation Process
Figure 2. Benchmark Sample Structure. A sample consists of an initial input followed by a sequence of instructions forming the prompt, which is processed by the model to generate the output.
Benchmark sample structure diagram.
@@
$3.$ How PACIFIC Meets the Criteria
@@ 5.1.1. Performance Across Difficulty Levels
Figure 3. Prompt Level Accuracy across varying complexity levels. Higher instruction counts and output lengths correspond to increased task difficulty. Average across all programming languages.
Prompt level accuracy graph
@@ 5.1.3. Model Comparison
Figure 4. Aggregate performance of evaluated LLMs, measured as the mean Prompt Level Accuracy across all benchmarks.
Aggregate performance of evaluated LLMs
@@
Figure 5. Aggregate missing answers rates across models and instruction counts. The higher the instruction count, the higher the error rate.
Aggregate missing answers rates
@@ 5.1.4. Instruction Samples
Figure 6. Accuracy across models for five representative instructions, grouped by complexity. The to_roman and next_perfect_squared are considered easy, the other ones are the hardest in our current instruction set. Scores are aggregated by instruction count and targeted output length.
Accuracy across models for five representative instructions
@@ A.1. Current Set of Natural Language Instructions
    Take the previous string - let’s call it s, and output the concatenation of s with itself two times $s+$s.
@@ A.3. Per-Model Performance Details
Figure 7. Claude-3.5-Haiku Performance. Results across all benchmarks for Metric Prompt Level Accuracy.
Claude-3.5-Haiku Performance
@@ A.3.2. Claude-4-Sonnet
Figure 8. Claude-4-Sonnet Performance. Results across all benchmarks for Metric Prompt Level Accuracy.
Claude-4-Sonnet Performance
@@ A.3.3. GPT-OSS-120B
Figure 9. GPT-OSS-120B Performance. Results across all benchmarks for Metric Prompt Level Accuracy.
GPT-OSS-120B Performance
@@ A.3.4. Llama-4-Maverick-17B-128E-Instruct
Figure 10. Llama-4-Maverick-17B-128E-Instruct Performance. Results across all benchmarks for Metric Prompt Level Accuracy.
Llama-4-Maverick-17B-128E-Instruct Performance
@@ A.3.5. Qwen3-235B-A22B-Instruct
Figure 11. Qwen3-235B-A22B-Instruct Performance. Results across all benchmarks for Metric Prompt Level Accuracy.
Qwen3-235B-A22B-Instruct Performance
@@ A.3.6. Qwen3-Coder-480B-A35B-Instruct
Figure 12. Qwen3-Coder-480B-A35B-Instruct Performance. Results across all benchmarks for Metric Prompt Level Accuracy.
Qwen3-Coder-480B-A35B-Instruct Performance